### v1

@@ Acknowledgements
ClinicalConcept, EvidenceStatement, Study, Guideline, Recommendation, Protocol, OutcomeMeasure, Device, AdverseEvent, Contraindication, Topic mentions, associated_with, belongs_to_topic, reported_in, supported_by, has_protocol, uses_device, reports_measure, recommends_for, recommends_against, has_adverse_event, contraindicated_for age_bin, age_range, sex, baseline_severity, time_since_injury_or_surgery, inclusion_criteria, exclusion_criteria, dose, frequency, session_duration, intensity, progression_rule, setting, supervision, followup_weeks, timepoint_value, timepoint_unit, measure_name, outcome_domain, primary_outcome, adverse_event, contraindication, recommendation_strength, evidence_certainty, applicability_notes, study_design, sample_size, protocol_params
S2. Prompt Templates
This section provides representative prompt excerpts used in SR-RAG. For brevity, we include only the core instructions and output constraints.
PICO-guided HyDE.
Listing S1: PICO-guided HyDE prompt excerpt.
You are a sports rehabilitation clinician.
- P: population characteristics and key constraints
- I: intervention details (dose/frequency/intensity/progression)
- C: comparator if applicable
- O: outcomes and relevant measures
- T: time horizon / follow-up
Constraints:
- Do NOT cite sources.
- Use neutral clinical language.
- 5-8 sentences, concise, declarative, paper-like.
Question:
{question}
Output ONLY the hypothetical evidence window.
In practice, we invoke this template multiple times per query to obtain diversified hypothetical windows.
Evidence window selection.
Listing S2: Evidence window selection prompt excerpt.
You are an evidence extraction assistant. Given a question and a medical passage,
identify up to MAX contiguous snippets from the passage that directly help answer the question.
Each snippet must be copied verbatim and kept short.
Return strictly JSON:
"windows": [
"snippet": "...",
"reason": "...",
"contains_population": true/false,
"contains_intervention": true/false,
"contains_comparator": true/false,
"contains_outcome": true/false,
"contains_time": true/false,
"outcome_and_time_same_window": true/false
Question: {question}
Passage: <<<{passage}>>>
Nugget coverage.
Listing S3: Nugget coverage prompt excerpt.
You are a clinical QA evaluator.
Given the question, a system answer, and ONE ground-truth nugget,
judge whether the nugget is explicitly covered by the system answer.
Rules:
- Only judge coverage by the system answer text itself.
- Treat synonymous expressions as covered.
- If the nugget is contradicted or missing, mark it as not covered.
- Output RFC8259-compliant JSON only.
Return JSON:
"covered": true/false,
"reason": "brief justification"
Question: {question}
System answer: {answer}
GT nugget: {nugget}
PICOT match accuracy.
Listing S4: PICOT match prompt excerpt.
You are a medical information extraction and matching assistant.
Given gold PICOT fields and the system answer, extract PICOT from the system answer and compare field-by-field.
Rules:
- Do not invent missing fields. If a field is not stated, output null.
- Consider medically equivalent synonyms as a match.
- Exclude GT fields that are explicitly marked as null/missing from scoring.
- Output RFC8259-compliant JSON only.
Return JSON:
"system_picot": {"P": "...|null", "I": "...|null", "C": "...|null", "O": "...|null", "T": "...|null"},
"match": {"P": 0/1, "I": 0/1, "C": 0/1, "O": 0/1, "T": 0/1},
"reason": {"P": "...", "I": "...", "C": "...", "O": "...", "T": "..."}
Question: {question}
Gold PICOT: {gold_picot}
System answer: {answer}
S3. Key Hyperparameters and BETR Training Details
Table S3 summarizes selected hyperparameters for SR-RAG and BETR. Evidence grades were defined as: A = guidelines/expert consensus; B = systematic reviews/meta-analyses; C = randomized controlled trials; D = cohort studies; E = other studies.
Table S3: Selected hyperparameters and training settings.
Item Setting
SR-RAG retrieval and reranking
Final evidence budget Top-$K$ contexts = 12; soft quota: retain $\geq$1 Grade A if its score is sufficiently high
Recall capacities dense top-$K$ = 300; graph top-$K$ = 120; HyDE passages retrieved via dense top-$K$ = 300; RRF $k$ = 60
HyDE 3 hypothetical passages per query; temperature = 0.3 for lexical diversification
Window selection window size = 320 tokens; max windows/chunk = 3; overlap = 64 tokens
Two-stage reranking ColBERT $\rightarrow$ cross-encoder
BETR
Data split query-disjoint train/validation: 1310/327 queries; 21,546/5,209 pairs
Pair construction 20 negatives per positive
Optimization Adam, learning rate 0.05, epochs 80, seed 42
Shrinkage strength $\tau$ selected via grid search on validation; selected $\tau=1.0$ and fixed for all experiments
Scale prior $\sigma_{a}=5.0$; reparameterized $a=\exp\{\alpha\}$
Learned parameters $a=1.0348$;
$u_{A}=0,\ u_{B}=-0.1287,\ u_{C}=-0.2575,\ u_{D}=-0.3863,\ u_{E}=-0.5151$
S4. Benchmark Composition by Sub-condition
Table S4 summarizes the SR-RAG benchmark composition (n = 1,637 queries). The benchmark covers 21 common sports rehabilitation sub-conditions, along with two cross-condition guideline sets. Specifically, the guideline sets include a general guideline set and a special-population subset.
Table S4: SR-RAG benchmark composition by sub-condition and guideline set.
Code Sub-condition / Guideline set Queries (n)
@@
MACL Knee meniscal and articular cartilage lesions 78

### v2

@@ C.3 Schema Tables
ClinicalConcept, EvidenceStatement, Study, Guideline, Recommendation, Protocol, OutcomeMeasure, Device, AdverseEvent, Contraindication, Topic mentions, associated_with, belongs_to_topic, reported_in, supported_by, has_protocol, uses_device, reports_measure, recommends_for, recommends_against, has_adverse_event, contraindicated_for participant_characteristics, procedure_details, measurement_details, temporal_details, adverse_event, contraindication, recommendation_strength, evidence_certainty, applicability_notes, study_design, sample_size, protocol_params
Note. The first four attributes are free-text fields that collapse the 20 PICO-specific attributes of Table C.1 into domain-generic buckets (participant_characteristics $\leftarrow$ P-related; procedure_details $\leftarrow$ I-related; measurement_details $\leftarrow$ O-related; temporal_details $\leftarrow$ T-related), preserving underlying clinical information while removing PICO-structured encoding.
Appendix D Pipeline Details and Benchmark Construction
D.1 Three-channel Retrieval Fusion
SR-RAG combines ranking results from three retrieval channels using Reciprocal Rank Fusion (RRF) (Cormack et al., 2009). For any candidate window $d$, the RRF score is defined as:
$\mathrm{RRF}(d)=\sum_{c\in\{\mathrm{Dense},\mathrm{Graph},\mathrm{HyDE}\}}\frac{1}{k+\mathrm{rank}_{c}(d)},$ (D.1)
D.2 HyDE Background
The original HyDE method addresses poor zero-shot dense retrieval performance without relevance annotations (Gao et al., 2023). LLMs generate several hypothetical documents related to a user query, which are subsequently used for dense retrieval in the corpus. This paradigm tolerates factual errors in hypothetical documents: because retrieval relies on dense embeddings rather than lexical matching, semantic similarity is preserved even when generated content contains inaccuracies (Gao et al., 2023). Subsequent work identified another advantage of HyDE: bridging the semantic gap between queries and documents (Wang et al., 2023). In RAG, document-form inputs generally outperform raw question-form queries; thus, HyDE effectively improves retrieval quality and downstream performance.
D.3 ColBERT MaxSim Mechanism
ColBERT uses the MaxSim mechanism for scoring:
$s_{\mathrm{col}}(q,d)=\sum_{i=1}^{|q|}\ \max_{j\leq|d|}\ \cos\!\left(\mathbf{e}_{i}^{q},\mathbf{e}_{j}^{d}\right)$ (D.2)
Here, $\mathbf{e}_{i}^{q}$ and $\mathbf{e}_{j}^{d}$ denote the $i$-th and $j$-th token embeddings of the query and window, respectively. The MaxSim mechanism computes, for each query token, its maximum similarity with any window token and sums across tokens (Khattab and Zaharia, 2020), improving sensitivity to medical abbreviations, scale names, and synonyms while increasing matching robustness.
D.4 Cross-encoder Details
After coarse ranking, we used BGE-reranker-v2-m3 as a cross-encoder to perform fine-grained ranking on the top-$K$ high-scoring candidates (Chen et al., 2024). The cross-encoder jointly encodes the query and each candidate window via concatenation, enabling full interaction through multiple self-attention layers and outputting an overall relevance logit $s(q,d)$. The computational cost of the cross-encoder increases linearly with the number of candidates. By adopting a two-stage pipeline, the system reduces reranking time and computational cost while balancing precise term matching and semantic understanding of long texts.
D.5 Benchmark Construction Full Pipeline
Stage 1: Question Generation.
Using the corpus with completed evidence-grade annotation and chunking, we performed stratified sampling by evidence grade to obtain candidate evidence windows, filtering those with insufficient information density or lacking substantive clinical conclusions. For each candidate window, GPT-4o (OpenAI, 2024) generated a clinical question strictly corresponding to the window’s core conclusion while simultaneously extracting PICOT elements. During generation, questions had to be directly supported by the window; inclusion of information not present in the window was prohibited (Saad-Falcon et al., 2024). Additionally, the model annotated evidence certainty (sufficient or uncertain) for each question, enabling stratified evaluation and error analysis.
Stage 2: Retrieval Accessibility Check.
We performed retrieval accessibility checks for each generated question, similar to the round-trip consistency filtering in ARES (Saad-Falcon et al., 2024). Using a system-agnostic hybrid retrieval baseline (combining sparse and dense methods), we retrieved the top-$K$ candidate windows. If the seed window was recalled within top-$K$, the question was deemed “accessible” and entered the main split; otherwise, it entered the challenge split for manual review.
Stage 3: Reference Evidence Construction.
We selected gold windows for each accessible question from a larger candidate pool. This entailed reranking candidates by relevance, then having DeepSeek-V3 (DeepSeek-AI, 2024) grade the evidence relationship (strongly supportive/supportive/weakly related/unrelated). Strongly supportive and supportive windows formed the gold evidence set. Based on gold windows, we generated hierarchical reference answers: an exact answer summarizing the core conclusion, and an ideal answer organized by evidence grades (A–E), with Grade A prioritized and remaining evidence presented in descending pyramid order. All statements were strictly aligned with gold windows, with source identifiers for traceability. GPT-4o then decomposed the exact answer into atomic facts (nuggets) for automated assessment (Min et al., 2023).
Stage 4: Human Review.
Four sports science graduate students manually reviewed all QA pairs from both the main and challenge splits, with particular attention to samples marked “evidence uncertain” and those in the challenge split. Questions that were clinically irrelevant, lacked practical rehabilitation value, or contained factual errors were discarded. Approximately 2,000 questions were generated initially; after screening, 1,637 were retained as the official evaluation set. We extracted questions, exact answers, and nuggets and compiled them into a public dataset for reusability. Because some gold windows originate from institutionally subscribed materials, the public version excluded evidence window text.
D.6 Benchmark Composition
Table D.1 summarizes the SR-RAG benchmark composition (n=1,637n=1{,}637 queries). The benchmark covers 21 common sports rehabilitation sub-conditions, along with two cross-condition guideline sets.
Table D.1: SR-RAG benchmark composition by sub-condition and guideline set.
Code Sub-condition / Guideline set $n$
@@ D.6 Benchmark Composition
MACL Meniscal and articular cartilage lesions 78
@@
Appendix E Evaluation Details
Nugget coverage and PICOT match accuracy were implemented via LLM-as-judge (Zheng et al., 2023): for each nugget, the LLM determined whether it was covered by the system answer; for each extracted P/I/C/O/T field, the LLM evaluated whether it matched the reference (paraphrasing permitted; null GT fields were excluded from scoring). Coverage and match rates were then aggregated. Answer faithfulness and semantic similarity were assessed via the RAGAS framework (Es et al., 2024). RAGAS first extracted claims from the answer, then determined for each claim whether it was supported by the actually retrieved context, yielding a faithfulness score (also LLM-as-judge based). Semantic similarity was computed as cosine similarity between text embeddings of the system answer and GT.
For manual evaluation, review materials included task instructions, scoring criteria, evaluation dimensions, and the question set. For each question, materials provided the question and GT, system-generated answer, retrieved evidence fragments, and a scoring sheet. Detailed scoring criteria for each dimension:
  • Medical factual accuracy: whether conclusions, values, and prescriptions align with current evidence, guidelines, and consensus.
  • Answer faithfulness: whether key statements are substantiated by retrieved fragments.
  • Answer relevance: whether the response addresses core concerns and covers key constraints.
  • Safety: whether the answer contains potentially harmful or misleading content.
  • PICOT alignment: whether the answer is structured around the question’s P/I/C/O/T elements without population or intervention mismatch.
We aggregated scores from all five experts across five dimensions for the 20 questions and computed means and standard deviations per question and dimension for analysis.
Appendix F Prompt Templates
Figure F.1 shows the PICO-guided HyDE prompt, the key generation template in SR-RAG.
You are a sports rehabilitation clinician.
- P: population characteristics and key constraints
- I: intervention details (dose/frequency/intensity/progression)
- C: comparator if applicable
- O: outcomes and relevant measures
- T: time horizon / follow-up
Constraints:
- Do NOT cite sources.
- Use neutral clinical language.
- 5-8 sentences, concise, declarative, paper-like.
Question:
{question}
Output ONLY the hypothetical evidence window.
Figure F.1: Prompt for PICO-guided HyDE generation.
Appendix G Implementation Details
All experiments were implemented in Python 3.11. Graph construction followed Youtu-GraphRAG’s schema-bounded extraction and community compression using NetworkX, with graph artifacts stored as JSON. Dense retrieval was implemented with sentence-transformers and FAISS; candidate lists from dense, graph, and HyDE retrieval were fused via RRF. Two-stage reranking used ColBERT (mxbai-edge-colbert-v0) for coarse ranking and a cross-encoder (bge-reranker-v2-m3) for final logits. BETR training used PyTorch 2.6.0 with the Adam optimizer (learning rate: 0.05, epochs: 80, $K=20$ negatives per positive). Hyperparameters ($\tau$) were selected via grid search on the validation split (nn=327). Unless otherwise specified, LLM decoding used temperature = 0.
Table G.1 summarizes selected hyperparameters for SR-RAG and BETR.
Table G.1: Selected hyperparameters and training settings.
Item Setting
SR-RAG retrieval and reranking
Final evidence budget Top-$K$ = 12
Recall capacities dense 300; graph 120; HyDE 300; RRF kk=60
HyDE 3 passages/query; temp = 0.3
Window selection 320 tokens; max 3/chunk; overlap 64
Two-stage reranking ColBERT $\rightarrow$ cross-encoder
BETR
Data split train/val/test: 983/327/327 queries
Pair construction 20 negatives per positive
Optimization Adam, lr 0.05, 80 epochs
Shrinkage $\tau$ grid search; selected $\tau$=1.0
Scale prior $\sigma_{a}$ 5.0
Appendix H Case Study
We present a representative case to illustrate how SR-RAG integrates EBM principles throughout the pipeline. The LLM used was DeepSeek-V3. Figure H.1 provides an overview of the case study workflow.
Figure H.1: Case-study workflow of SR-RAG, illustrating the full pipeline from query to structured output.
Query.
“In cardiopulmonary rehabilitation programs, under what circumstances should clinicians choose high-intensity interval training (HIIT) over moderate-intensity continuous training (MICT) for prepubertal patients with congenital heart disease?” This question exhibits a clear PICO structure and requires cross-evidence integration and conditional reasoning.
PICO-guided HyDE.
The model generated three hypothetical documents with PICO soft constraints (Figure H.1), converting the query into declarative hypothetical answers. The model preferentially reused extractable P/I/C/O/T anchors from the query, tolerating missing fields but prohibiting hallucination of values for absent fields. This bridges the semantic gap between the query and ground truth evidence.
EBM-adaptive Retrieval.
SR-RAG ran separate candidate generation on the Grade A corpus and Grades B–E corpus with their own recall quotas, then merged the candidate sets. BETR calibration was applied to the merged candidates, ensuring that while semantic relevance remains dominant, higher grades were prioritized when relevance is comparable.
Three-channel Fusion and Iterative Reasoning.
Three retrieval channels (Dense, Graph, HyDE) were fused via RRF. The IRCoT module performed three iterations of schema-guided reasoning and reflection (Figure H.1).
Two-stage Reranking and Selection.
ColBERT applied the MaxSim mechanism for coarse ranking; the cross-encoder performed fine-grained ranking within top-$K$. BETR calibration produced the final ranking score $r(q,d)=\hat{a}\,s(q,d)+\hat{u}_{\mathrm{Grade}(d)}$. In this case, two key guideline windows entered the top-2, replacing candidates with potential population mismatch risks (Figure H.1).
Structured Output.
The first paragraph concisely summarizes the core answer; the second elaborates on guidelines and supporting evidence; the third compares the query and response, highlighting limitations (Figure H.1).
Observations.
Three key findings emerge from this case: (i) PICO-guided HyDE successfully extracted anchors (Population: prepubertal CHD patients; Intervention: HIIT; Comparator: MICT; Outcome: cardiopulmonary function) without fabricating missing fields, reducing retrieval drift. (ii) The dual-track retrieval strategy successfully recalled relevant guideline evidence that would have been diluted in a single-pool retrieval, demonstrating the value of evidence-grade-aware candidate generation. (iii) BETR calibration promoted a Grade A guideline window over a Grade C RCT with slightly higher semantic similarity, aligning the final ranking with EBM principles.